OpenStudy (anonymous):

Help!!! 3(5)^2x-3 =6

OpenStudy (anonymous):

\[3\times 5^{2x}-3=6\] \[3\times 5^{2x}=9\] \[5^{2x}=3\] is a start

OpenStudy (anonymous):

unless i read it wrong is it \[3\times 5^{2x-3}=9\]?

OpenStudy (anonymous):

if it is the latter, then start with \[5^{2x-3}=3\] which one might it be?

OpenStudy (anonymous):

The first one is correct the one you started

OpenStudy (anonymous):

Actually it's the latter

OpenStudy (anonymous):

ok so now it looks like we want it \[5^{2x}=3\] then\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] you have \[5^{2x}=3\] so \[2x=\frac{\ln(3)}{\ln(5)}\] making \[x=\frac{\ln(3)}{2\ln(5)}\]

OpenStudy (anonymous):

Except it equals6

OpenStudy (anonymous):

I'm sorry

OpenStudy (anonymous):

is it this \[3\times 5^{2x-3}=9\]

OpenStudy (anonymous):

Yeah but it equals 6 instead of 9

OpenStudy (anonymous):

\[3\times 5^{2x-3}=6\]?

OpenStudy (anonymous):

divide by 3, get \[5^{2x-3}=2\]

OpenStudy (anonymous):

then \[2x-3=\frac{\ln(2)}{\ln(5)}\]

OpenStudy (anonymous):

add three and divide by 2 to solve for \(x\)

OpenStudy (anonymous):

Okay Is the answer x=1.71